OpenStudy (anonymous):

why does this integral not exist ∫-5 to 5 (10/(x^3))dx ? and this one ∫x to 2x -10csc^2(x) dx ?

OpenStudy (johnweldon1993):

Not exist? Well it does exist... \[\large \int_{-5}^{5} \frac{10}{x^3}dx\] \[\large 10\int_{-5}^{5} \frac{1}{x^3}dx\] \[\large 10\int_{-5}^{5} x^{-3}dx\] if you do that out...you do get a number...what do you get?

OpenStudy (anonymous):

well the computer says that's the correct answer so idk

OpenStudy (amistre64):

how can you integrate across a domain that includes a value thats not in the domain?

OpenStudy (amistre64):

0 is a bad value for dividing by

OpenStudy (amistre64):

\[\int_{-5}^{5}10x^{-3}~dx=\lim_{t\to 0}\left(\int_{-5}^{t}10x^{-3}~dx+\int_{t}^{5}10x^{-3}~dx\right)\]

OpenStudy (amistre64):

if the limit exists, then the integration is defined by it; if it doesnt exist, then the integration does not converge

OpenStudy (anonymous):

thanks!